OpenStudy (anonymous):

Help please, If the point P in the unit circle that corresponds to a real number t is [5/7, -sqrt6:radicand2/7] find csc(t)?

OpenStudy (anonymous):

The expression is a bit confusing. Can you rewrite that?

OpenStudy (anonymous):

Did you mean \[(\frac{5}{7}, \frac{-2\sqrt6}{7})\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am having so much trouble learning trig at the age of 45 lol

OpenStudy (anonymous):

Any point on the unit circle can be represented by \[(\sin(t), \cos(t))\]. The angle t isn't important here. What this tells us is that the current angle has a sine of (5/7). And we know that csc(t) = 1/sin(t), so all we do is flip the sine value to get csc(t) for the t in question = 7/5

OpenStudy (anonymous):

oh my.........thank you so much

OpenStudy (anonymous):

so I need to study the trig functions of sin cos, and tan, and then flip for cot, csc, and sec

OpenStudy (anonymous):

You have the right name, lifesaver!

OpenStudy (anonymous):

Make sure you know which one is the inverse of which: csc(t) = 1/sin(t) sec(t) = 1/cos(t) cot(t) = 1/tan(t) just in case

OpenStudy (anonymous):

So......if I have this P on this unit circle that corresponds to a real number (t) and the [-sqrt:7/4, -3/4] and I need to find cot(t) the answer would be the square root of 7/3 right?

OpenStudy (anonymous):

Depends on where you mean to have parenthesis, but even then, still not quite: (sqrt(7))/3 = tan(t). So cot(t) is the inverse of that: 3/(sqrt(7))